# Enhancing Weakly Supervised 3D Medical Image Segmentation through Probabilistic-aware Learning

### v1

@@ Abstract
Keywords:
3D medical imaging weakly supervised learning segmentation probabilistic-based framework
@@ 2 Related Work
2.0.1 Medical Image Segmentation.
@@
2.0.2 Weakly Supervised Segmentation.
Weakly supervised learning reduces annotation cost by using sparse annotations instead of fully annotated masks. Weak labels such as bounding boxes [32, 8], scribbles [21], and points [2] have been utilized. Zhang et al. [37] integrates point-level annotation and sequential patch learning for CT segmentation. Roth et al. [31] designs a point-based loss function with an attention mechanism. Zou et al. [40] proposes a well-calibrated pseudo-labeling strategy, while Liu et al. [22] introduces an informative selection strategy. In contrast, our work proposes a “dense" weak annotation approach from a probabilistic perspective.
2.0.3 Probabilistic Modeling in Deep Learning.
@@
Figure 1: Overview of our framework. We adopt UNETR[14] as the baseline network for our segmentation model. The input is a 3D medical volume, which is processed by our Probabilistic Transformer Network, which is powered by the PMSA mechanism. The output of the network is a 3D segmentation map, which is supervised by the “dense" probability-based pseudo label generated from “sparse" point-based annotations. A Probability-aware Segmentation Loss Function is proposed to train the network.
@@ 3.1 Overview
Medical image segmentation, which is typically referred to as semantic segmentation of medical images, aims to partition the image into different non-overlapping regions with unique semantic labels. Given an image $I$ and the semantic classes $\{C_{1},C_{2},...,C_{k}\}$, the semantic segmentation process is performed by dividing $I$ into $\{D_{1},D_{2},...,D_{k}\}$ (i.e. the subregions), which satisfies:
@@
Fig. 1 illustrates the overview of our method for solving the weakly supervised 3D medical image segmentation task. we introduce a novel weakly supervised training pipeline for 3D medical image segmentation, taking probabilistic features of both annotation process and network training into consideration. We illustrate our pipeline in the following aspects: 1) A probability-based pseudo label generation scheme for generating “dense" weak annotations. 2) A probabilistic Transformer network, whose key component is the proposed gaussian-based multi-head self-attention mechanism. 3) The probability-informed loss function.
@@ 3.2 Probability-based Pseudo Label Generation
3.2.1 Sparse Labels Annotation.
@@
3.2.2 Pseudo Label Generation.
@@
This process, applied to all annotated points, generates dense 3D labels by summing up the label maps from all points and normalizing the intensity to $[0,1]$. This probability-based pseudo label generation scheme effectively transforms sparse annotations into informative dense labels, improving the training of the segmentation network. The entire probability-based pseudo Label generation pipeline is illustrated in Fig. 2 and Appendix 0.A.
@@ 3.3 Probabilistic Transformer Network
Though the proposed pseudo label in Sec. 3.2.2 can reflect the confidence level of the annotator, the within-class variance is high, illustrated in Fig. 3, due to: 1) the inherent morphological variation of a human organ, 2) the randomness of the point-sampling process. Therefore, a probabilistic model is expected to model the complex distribution.
Another important feature of the proposed pseudo label is that the confidence of a specific point has latent correlations with the confidence of its surroundings. Considering that Vision-Transformer-based architecture can capture the long-range dependencies and global context of images, we introduce a probabilistic transformer network.
Figure 3: The ground truth label and the corresponding probability-based pseudo label of the spleen. The three pairs of images belong to different subjects, and we visualize the middle slice of the foreground of the ground truth label in the axial plane. The within-class variance of the pseudo label is influenced by both the inherent morphological variation of human organs as well as the randomness of the point-sampling process.
3.3.1 Network Architecture.
Our framework adopts the contracting-expanding schema characteristic of the UNETR architecture. Initially, a 3D volume $x\in\mathbb{R}^{H\times W\times D\times C}$, with dimensions $(H,W,D)$ and $C$ input channels, is segmented into non-overlapping uniform patches of dimensions $(P,P,P)$. This segmentation transforms the volume into a sequence $x_{v}\in\mathbb{R}^{N\times(P^{3}\cdot C)}$ by flattening these patches, where $N=(H\times W\times D)/P^{3}$ denotes the sequence length. Subsequently, these patches are mapped into a $K$-dimensional embedding space via a linear layer. Furthermore, a 1D learnable positional embedding $E_{pos}\in\mathbb{R}^{N\times K}$ is incorporated into the mapped patches. The process can be defined as follows:
@@
The output of the Probabilistic Transformer blocks, denoted as $z_{i}$ (where $i$ takes values 3, 6, 9, 12), has a shape of $\frac{H\times W\times D}{P^{3}}\times K$ and is reshaped into $\frac{H}{P}\times\frac{W}{P}\times\frac{D}{P}\times K$.
In the decoder network, each feature $z_{i}$ undergoes deconvolution blocks to increase the resolution by a specific factor (2 for $z_{12}$ and $z_{9}$, 4 for $z_{6}$, and 8 for $z_{3}$). Starting from the lowest resolution, i.e., $z_{12}$ and $z_{9}$, the features are concatenated and unsampled to match the resolution of higher-level features. This procedure is iterated until the original resolution is fully reinstated. Subsequently, the output layer utilizes the feature map with the complete resolution to predict the final segmentation results. Within this network, the most critical component is PMSA, which we will delineate in the following part.
@@ 3.3.2 Probabilistic Multi-head Self-Attention
However, the Probability-based Pseudo Label suffers from large in-class variance caused by the randomness of the point-sampling process and the inherent diversity of human organ structure. To guide our model to capture the variance within the proposed pseudo label and encode the input properly, inspired by  [12], we introduce our Probabilistic Multi-head Self-Attention module. In a single SA head, we assume that the dependency score $\alpha_{ij}$ follows a Gaussian distribution: $\alpha_{ij}\sim\mathcal{N}(\mu_{ij},\sigma^{2}_{ij})$, where the mean $\mu_{ij}$ and the variance $\sigma^{2}_{ij}$ are calculated with $q_{i}$ and $k_{j}$ using a multilayer perceptron (MLP). In order to allow the parameters to be updated through backpropagation, we adopt reparameterization trick [18]:
$\rm\alpha_{ij}=\mu_{ij}+\sigma_{ij}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1)$ (7)
@@
where $\alpha_{m}$ denotes the dependency scores sampled at the $m$th time, and $y^{*}$ is the final segmentation output. More details about the sampling of dependency scores and the proof of Eq. 9 can be found in Appendix 0.B.
3.4 Probability-informed Segmentation Loss Function.
As discussed in Sec. 3.2, the proposed pseudo label is considered a probability map, where the intensity of each point represents the annotator’s confidence in classifying it as the target organ. Therefore, to enable our model to be aware of the underlying confidence within the pseudo label, we introduce a loss function which is a combination of DICE loss and Probability-weighted Cross Entropy (PCE) loss. The intuition is that points with prior confidence greater than a certain threshold are considered as the foreground of the basic label map, while we weight the loss function with the prior confidence of the annotator since voxels with low confidence deserve lower loss weights.
@@
$\mathcal{L}_{\text{PCE}}=\frac{1}{N}\sum{i=1}^{N}\mathcal{L}_{\text{pce}}(p_{i},s_{i})$ (13)
@@
Here, $l$ and $h$ denote the index of the transformer block and head, respectively. $\alpha_{lhij}$ represents the dependency score sampled from the distribution $\mathcal{N}(\mu_{lhij},\sigma_{lhij}^{2})$, while $\alpha_{lhij}^{\prime}$ is calculated as the scaled dot-product of $q_{lhi}$ and $k_{lhj}$. By minimizing the KL loss, we encourage the distribution $\mathcal{N}(\mu_{lhij},\sigma_{lhij}^{2})$ to closely match $\mathcal{N}(\mu_{lhij}^{\prime},\sigma^{2})$, where $\sigma$ is empirically set to 1.
@@
$\ \mathcal{L}_{total}=\mathcal{L}_{Seg}+w\mathcal{L}_{KL}.$ (15)
@@ 4.1 Implementation Details
We conducted experiments on two authoritative datasets in the fields of CT and MRI, respectively: the BTCV dataset [19] — a part of the MICCAI 2015 Challenge, comprising multi-organ abdominal 3D CT scans acquired during the portal venous contrast phase; and the CHAOS dataset [17] — involving the segmentation of four abdominal organs from MRI datasets acquired with two different sequences (T1-DUAL and T2-SPIR). We assessed the effectiveness of our methodology using two prevalent metrics: the DICE score, where higher scores indicate better performance, and the 95% Hausdorff Distance (HD95), where lower values are preferable. For further details on the implementation, training, and inference strategies, please consult Appendix 0.C and Appendix 0.D.
@@ 4.2 Results
Comparison with state-of-the-art pseudo label generation methods. Tab. 1 shows the quantitative results for four organs: spleen, liver, left kidney, and right kidney. We categorize the weakly supervised methods into two types of supervision: point-supervised learning and scribble-supervised learning. Point-supervised methods use a few annotated points to guide the segmentation, such as sparse [7], convex [1] and ours. ADNet [13] and ALPNet [26] are examples of using scribbles-supervised learning to generate pseudo labels.
Table 1: Comparison with SOTA weakly supervised methods on the BTCV and CHAOS datasets using the Dice metric.
Dataset Method Spleen Liver Left Kidney Right Kidney
@@
Comparison with SOTA fully supervised methods. To underscore the efficacy of our proposed approach, we juxtapose our weakly supervised method against state-of-the-art fully supervised methods in BTCV dataset, including TransUnet [5], SwinUnet [3], UCTransNet [35] and UNETR[14]. It is paramount to note that this juxtaposition is inherently imbalanced, as our method operates on notably sparser original annotations compared to the comprehensive annotations utilized by the aforementioned fully supervised methods.
Despite this inherent disparity, as delineated in Tab. 2, our method exhibits performances that are remarkably on par with, and in certain metrics, even surpass, those achieved by fully supervised counterparts. For instance, our method eclipses UCTransNet in spleen segmentation, showcasing the distinct advantages of our probabilistic weakly supervised approach. We present two visual illustrations of experimental results for four organs and their segmentation predictions in Appendix 0.E.
Figure 4: Qualitative comparison with weakly supervised methods.
@@
Method Spleen Liver Left Kidney Right Kidney
@@
Weakly Ours 0.8279 63.09 0.8157 265.79 0.7599 266.17 0.7164 116.22
@@ 4.3 Ablation Study
Spleen Liver Left Kidney Right Kidney Spleen Liver Left Kidney Right Kidney
@@
Table 4: Quantitative results of ablation study for network structure and loss function. For these parts of comparison, only the methods corresponding to that part are varied, while the methods in the other parts are kept constant. “Ours” column represents the results of our complete approach.
@@
Left Kidney 84.26 106.01 133.96 106.79 46.39 31.59 93.48
Right Kidney 114.14 98.64 168.17 57.92 37.75 56.69 36.01
Effectiveness of Probability-based Pseudo Label Generation. We assess the performance of a probabilistic mechanism using a semi-supervised technique with random point selection at various thresholds, aiming to mimic real-world scenarios with irregular point distributions or feature absence. Our goal is to ascertain if this approach yields consistent results across typical real-world data distributions, bridging the gap between laboratory and real-life settings, and ensuring effectiveness in both controlled and varied authentic environments.
Effectiveness of Probabilistic Transformer Network Structure. We investigate the impact of the probabilistic mechanism in the network architecture. Tab. 4 presents our experimental results, comparing the performance of the Self-Attention (SA) and the Multi-head Self-Attention (MSA) methods. These results indicate the heightened accuracy and reliability of PMSA in producing segmentation results that closely align with the actual anatomical structures, demonstrating the significance of considering probabilistic modeling in our transformer network.
Effectiveness of Probability-informed Segmentation Loss Function. We examine the effectiveness of our designed loss function, as presented in Tab. 4. The conclusive results underscore the outstanding efficacy of our approach. Our approach consistently achieves higher effectiveness scores, demonstrating its ability to deliver more accurate and coherent segments. Compared to our approach, the existing non-probabilistic loss functions, specifically DICE, Cross-Entropy (CE), combined Dice-Cross-Entropy (DCE), and Focal demonstrate suboptimal performance, especially in segmenting the liver and both kidneys. These findings underscore the limitations of the existing loss functions and underscore the superiority of our designed probability-informed loss function in achieving improved 3D medical image segmentation results.
Exploring Key Parameters —— Number of Sampled Points and Selection of Variance. In our study, we conduct in-depth ablation analyses on two crucial parameters. Specifically, the number of sampled points, as detailed in our annotation strategy (Sec. 3.2), plays a pivotal role in pseudo label generation. Additionally, the selection of variance $\sigma^{2}$ in computing the KL loss, a critical hyperparameter, is meticulously evaluated to determine its influence on segmentation accuracy. Detailed comparative experiments for both key parameters are conducted across two datasets, with the results comprehensively documented in two tables available in Appendix 0.F.

### v2

@@ Abstract
[…]
[cmu]organization=Carnegie Mellon University, city=Pittsburgh, postcode=15213, state=PA, country=USA
[…]
[towson]organization=Towson University, city=Towson, postcode=21252, state=MD, country=USA
[…]
[uab]organization=University of Alabama at Birmingham, city=Birmingham, postcode=35294, state=AL, country=USA
[…]
[umd]organization=University of Maryland, city=College Park, postcode=20742, state=MD, country=USA
@@ 2 Related Work
2.1 Medical Image Segmentation
@@
2.2 Weakly Supervised Segmentation
Weakly supervised learning reduces annotation cost by using sparse annotations instead of fully annotated masks. Weak labels such as bounding boxes [23, 24], scribbles [17], and points [16] have been utilized. Zhang et al. [25] integrates point-level annotation and sequential patch learning for CT segmentation. Roth et al. [26] designs a point-based loss function with an attention mechanism. Zou et al. [27] proposes a well-calibrated pseudo-labeling strategy, while Liu et al. [28] introduces an informative selection strategy. In contrast, our work proposes a ”dense” weak annotation approach from a probabilistic perspective.
2.3 Probabilistic Modeling in Deep Learning
@@
Figure 1: Overview of our framework. We adopt UNETR [33] as the baseline network for our segmentation model. The input is a 3D medical volume, which is processed by our Probabilistic Transformer Network, which is powered by the PMSA mechanism. The output of the network is a 3D segmentation map, which is supervised by the “dense” probability-based pseudo label generated from “sparse” point-based annotations. A Probability-aware Segmentation Loss Function is proposed to train the network.
@@ 3.1 Overview
Medical image segmentation, which is typically referred to as semantic segmentation of medical images, aims to partition the image into different non-overlapping regions with unique semantic labels. Given an image $I$ and the semantic classes $\{C_{1},C_{2},...,C_{k}\}$, the semantic segmentation process is performed by dividing $I$ into $\{D_{1},D_{2},...,D_{k}\}$ (i.e., the subregions), which satisfies:
@@
Fig. 1 illustrates the overview of our method for solving the weakly supervised 3D medical image segmentation task. We introduce a novel weakly supervised training pipeline for 3D medical image segmentation, taking probabilistic features of both annotation process and network training into consideration. We illustrate our pipeline in the following aspects: 1) A probability-based pseudo label generation scheme for generating “dense” weak annotations. 2) A probabilistic Transformer network, whose key component is the proposed gaussian-based multi-head self-attention mechanism. 3) The probability-informed loss function.
@@ 3.2 Probability-based Pseudo Label Generation
3.2.1 Sparse Labels Annotation
@@
3.2.2 Pseudo Label Generation
@@
This process, applied to all annotated points, generates dense 3D labels by summing up the label maps from all points and normalizing the intensity to $[0,1]$. This probability-based pseudo label generation scheme effectively transforms sparse annotations into informative dense labels, improving the training of the segmentation network. The entire Probability-based Pseudo Label Generation pipeline is illustrated in Fig. 2, while the algorithm of target generation is described in Algorithm 1.
Input: $C=\{x_{i},y_{i},z_{i}\}_{i=1}^{n}$: coordinates of sampled points, $X$, $Y$, $Z$
Output: $P\in{\mathbb{R}^{X\times{Y}\times{Z}}}$: label map
1 for $\{x_{i},y_{i},z_{i}\}$ in $C$ do
2       $P_{i}\leftarrow{0}$;
3       for $x\leftarrow{0}$ to $X$ do
4             for $y\leftarrow{0}$ to $Y$ do
5                   for $z\leftarrow{0}$ to $Z$ do
6                         $P_{i}(x,y,z)=e^{-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}+(z-z_{i})^{2}}{2\sigma^{2}}}$;
7
8
9
10
[…]
12 Normalize $P$ to $[0,1]$;
Algorithm 1 Target Generation
@@ 3.3 Probabilistic Transformer Network
Though the proposed pseudo label in Section 3.2.2 can reflect the confidence level of the annotator, the within-class variance is high, due to the inherent morphological variation of human organs and the randomness of the point-sampling process. Therefore, a probabilistic model is expected to capture the complex distribution.
3.3.1 Network Architecture
Our framework adopts the contracting-expanding schema characteristic of the UNETR architecture, as illustrated in Fig. 3. Initially, a 3D volume $x\in\mathbb{R}^{H\times W\times D\times C}$, with dimensions $(H,W,D)$ and $C$ input channels, is segmented into non-overlapping uniform patches of dimensions $(P,P,P)$. This segmentation transforms the volume into a sequence $x_{v}\in\mathbb{R}^{N\times(P^{3}\cdot C)}$ by flattening these patches, where $N=(H\times W\times D)/P^{3}$ denotes the sequence length. Subsequently, these patches are mapped into a $K$-dimensional embedding space via a linear layer. Furthermore, a 1D learnable positional embedding $E_{pos}\in\mathbb{R}^{N\times K}$ is incorporated into the mapped patches. The process can be defined as follows:
@@ 3.3.1 Network Architecture
Figure 3: Network Architecture Overview
@@ 3.3.2 Probabilistic Multi-head Self-Attention
However, the Probability-based Pseudo Label suffers from large in-class variance caused by the randomness of the point-sampling process and the inherent diversity of human organ structure. To guide our model to capture the variance within the proposed pseudo label and encode the input properly, inspired by Guo et al. [32], we introduce our Probabilistic Multi-head Self-Attention module. In a single SA head, we assume that the dependency score $\alpha_{ij}$ follows a Gaussian distribution: $\alpha_{ij}\sim\mathcal{N}(\mu_{ij},\sigma^{2}_{ij})$, where the mean $\mu_{ij}$ and the variance $\sigma^{2}_{ij}$ are calculated with $q_{i}$ and $k_{j}$ using a multilayer perceptron (MLP). In order to allow the parameters to be updated through backpropagation, we adopt reparameterization trick [34]:
$\alpha_{ij}=\mu_{ij}+\sigma_{ij}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1)$ (7)
@@
where $\alpha_{m}$ denotes the dependency scores sampled at the $m$th time, and $y^{*}$ is the final segmentation output. More details about the sampling of dependency scores and the proof of Eq. 9 can be found in  A.
3.4 Probability-informed Segmentation Loss Function
As discussed in Section 3.2, the proposed pseudo label is considered a probability map, where the intensity of each point represents the annotator’s confidence in classifying it as the target organ. Therefore, to enable our model to be aware of the underlying confidence within the pseudo label, we introduce a loss function which is a combination of DICE loss and Probability-weighted Cross Entropy (PCE) loss. The intuition is that points with prior confidence greater than a certain threshold are considered as the foreground of the basic label map, while we weight the loss function with the prior confidence of the annotator since voxels with low confidence deserve lower loss weights.
@@ 3.4 Probability-informed Segmentation Loss Function
$\mathcal{L}_{\text{PCE}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\text{pce}}(p_{i},s_{i})$ (13)
@@
$\mathcal{L}_{total}=\mathcal{L}_{Seg}+w\mathcal{L}_{KL}$ (15)
@@
3.5 Training and Inference Strategy
Binary Classification
Unlike binary segmentation map which is common in most deep learning tasks, the proposed probability-based pseudo label suffers from large data size and a single point has confidence scores for multiple organ classes, the sum of which might be greater than 1, which could be ambiguous. Thus, our model is trained to make inference of a single organ class, which is formulated as a binary classification task for each point.
Sampling of Dependency Scores
During training, to accelerate the training process, we sample the dependency scores for only one time, while during inference, the dependency scores are sampled for $M$ times, and the final output is calculated as Eq. 10.
@@ 4.1 Implementation Details
Our model inherits the contracting-expanding pattern of UNETR [33] but substitutes the encoder with a stack of Probabilistic Transformer blocks, each connected to the decoder through skip connections. We implemented our method using PyTorch [35] and MONAI [^1] https://monai.io/. All experiments were conducted on a single NVIDIA RTXA5000 GPU with 24GB of memory. We set the number of transformer encoders to 12 (L=12) with an embedding size of 768 (K=768). Each patch has a resolution of 16x16x16.
During training, we used the AdamW optimizer with an initial learning rate of 0.0001 and a batch size of 1, over 6,000 iterations. For inference, we employed a sliding window approach with a 50% overlap. The number of sampled points for different labels is proportional to the volume of the corresponding organ: 200 points for the spleen, 400 points for the liver, and 50 points for each of the right and left kidneys.
We conducted experiments on two authoritative datasets: the BTCV dataset [36], which includes multi-organ abdominal 3D CT scans acquired during the portal venous contrast phase, and the CHAOS dataset [37], involving the segmentation of four abdominal organs from MRI datasets acquired with two different sequences (T1-DUAL and T2-SPIR). We assessed the effectiveness of our methodology using two prevalent metrics: the DICE score, where higher scores indicate better performance, and the 95% Hausdorff Distance (HD95), where lower values are preferable.
@@ 4.2 Results
Comparison with state-of-the-art pseudo label generation methods
Tab. 1 shows the quantitative results for four organs: spleen, liver, left kidney, and right kidney. We categorize the weakly supervised methods into two types of supervision: point-supervised learning and scribble-supervised learning. Point-supervised methods use a few annotated points to guide the segmentation, such as sparse [22], convex [38] and ours. ADNet [39] and ALPNet [40] are examples of using scribbles-supervised learning to generate pseudo labels.
Table 1: Comparison with SOTA weakly supervised methods on the BTCV and CHAOS datasets using the Dice metric. ”LK” stands for ”Left Kidney”, and ”RK” stands for ”Right Kidney”. Throughout this text, these abbreviations will consistently be used to refer to these anatomical structures.
Dataset Method Spleen Liver LK RK
@@
Comparison with SOTA fully supervised methods
To highlight the efficacy of our proposed approach, we compare our weakly supervised method with state-of-the-art fully supervised methods on the BTCV dataset, including TransUnet [1], SwinUnet [6], UCTransNet [5] and UNETR [33]. It is paramount to note that this comparison is inherently imbalanced, as our method operates on notably sparser original annotations compared to the comprehensive annotations utilized by the aforementioned fully supervised methods.
@@
Method Spleen Liver LK RK
@@
Weakly Ours 0.8279 63.09 0.8157 127.16 0.7599 135.88 0.7164 116.22
Figure 4: Qualitative comparison with weakly supervised methods.
Despite this inherent disparity, as delineated in Tab. 2, our method exhibits performances that are remarkably on par with, and in certain metrics, even surpass, those achieved by fully supervised counterparts. For instance, our method eclipses UCTransNet in spleen segmentation, showcasing the distinct advantages of our probabilistic weakly supervised approach.
To provide a more intuitive understanding of our method’s performance, we present qualitative segmentation results in Fig. 5, which demonstrates our model’s capability to accurately delineate organ boundaries across different anatomical structures. The visual results confirm that our weakly supervised approach can achieve segmentation quality comparable to fully supervised methods, while requiring significantly less annotation effort.
Figure 5: Qualitative results of segmentation prediction.
In Fig. 6, we provide visual comparisons of our method against several fully supervised approaches, demonstrating its effectiveness in accurately segmenting key regions of interest. Despite limited supervision during training, our approach achieves performance comparable to fully supervised methods, underscoring its potential as a strong alternative.
Figure 6: Qualitative comparison with fully supervised methods.
@@ 4.3 Ablation Study
Effectiveness of Probability-based Pseudo Label Generation
We assess the performance of a probabilistic mechanism using a semi-supervised technique with random point selection at various thresholds, aiming to mimic real-world scenarios with irregular point distributions or feature absence. Our goal is to ascertain if this approach yields consistent results across typical real-world data distributions, bridging the gap between laboratory and real-life settings, and ensuring effectiveness in both controlled and varied authentic environments.
@@
Spleen Liver LK RK Spleen Liver LK RK
@@
Figure 7: Visualization results on BTCV dataset showing the comparison of pseudo-labels generated under different thresholds of random original labels.
Figure 8: Visualization results on CHAOS dataset demonstrating the effectiveness of our method across different threshold settings.
Fig. 7 and Fig. 8 show a comparison of the visualizations of pseudo-labels generated by our method under different thresholds of random original labels in two multi-modal datasets. Since our method is based on sparse label generation, the pseudo-labels produced by our approach tend to be much smoother compared to the ground truth. This is a shortcoming of our method, as well as a challenge that needs to be addressed in the field of weak supervision. Consequently, the boundary information is slightly inferior, which also explains the slightly higher HD95 distance. However, our method can still produce pseudo-labels with an overall contour even when sampling from randomly labeled points with very low thresholds. This makes our method more advantageous for datasets with very few or poorly labeled annotations.
Table 4: Quantitative results of ablation study for network structure and loss function. For these parts of comparison, only the methods corresponding to that part are varied, while the methods in the other parts are kept constant. ”Ours” column represents the results of our complete approach.
@@
LK 84.26 106.01 133.96 106.79 46.39 31.59 93.48
RK 114.14 98.64 168.17 57.92 37.75 56.69 36.01
Effectiveness of Probabilistic Transformer Network Structure
We investigate the impact of the probabilistic mechanism in the network architecture. Tab. 4 presents our experimental results, comparing the performance of the Self-Attention (SA) and the Multi-head Self-Attention (MSA) methods. These results indicate the heightened accuracy and reliability of PMSA in producing segmentation results that closely align with the actual anatomical structures, demonstrating the significance of considering probabilistic modeling in our transformer network.
Effectiveness of Probability-informed Segmentation Loss Function
We examine the effectiveness of our designed loss function, as presented in Tab. 4. The conclusive results underscore the outstanding efficacy of our approach. Our approach consistently achieves higher effectiveness scores, demonstrating its ability to deliver more accurate and coherent segments. Compared to our approach, the existing non-probabilistic loss functions, specifically DICE, Cross-Entropy (CE), combined Dice-Cross-Entropy (DCE), and Focal demonstrate suboptimal performance, especially in segmenting the liver and both kidneys. These findings underscore the limitations of the existing loss functions and underscore the superiority of our designed probability-informed loss function in achieving improved 3D medical image segmentation results.
4.4 Parameters Exploration
In our study, we conduct in-depth ablation analyses on two crucial parameters. Specifically, the number of sampled points, as detailed in our annotation strategy (Section 3.2), plays a pivotal role in pseudo label generation. Additionally, the selection of variance $\sigma^{2}$ in computing the KL loss, a critical hyperparameter, is meticulously evaluated to determine its influence on segmentation accuracy.
Table 5: Illustration of the number of sampled points.
Dataset n Dice↑ HD95↓
Spleen Liver LK RK Spleen Liver LK RK
BTCV 50 0.6392 0.1081 0.6276 0.5678 336.25 127.16 135.88 292.41
100 0.8001 0.7307 0.6366 0.4772 174.41 295.54 197.66 312.45
150 0.5462 0.7164 0.3856 0.5756 348.11 304.69 333.50 183.65
200 0.8279 0.8157 0.7599 0.7164 63.08 265.79 266.17 116.22
CHAOS 200 0.7356 0.8205 0.6662 0.6530 77.31 74.05 93.48 94.50
250 0.6982 0.7698 0.5687 0.7716 148.26 103.01 107.56 36.01
300 0.7402 0.7970 0.5978 0.6667 130.12 48.79 103.09 112.75
350 0.6711 0.7872 0.5516 0.6627 53.11 48.75 101.34 99.07
Impact of the Number of Sampled Points
Table 6: Illustration of the number of selection of variance.
Dataset $\sigma^{2}$ Dice Score↑ HD95↓
Spleen Liver LK RK Spleen Liver LK RK
BTCV 0.1 0.5104 0.7478 0.5474 0.5192 394.87 299.2 329.39 273.61
1 0.8279 0.8157 0.7599 0.7164 63.09 265.79 266.17 116.22
10 0.5754 0.7691 0.6103 0.3468 388.11 312.36 238.93 323.79
CHAOS 0.1 0.5472 0.5707 0.3669 0.3813 56.26 51.06 31.72 56.14
1 0.7402 0.8205 0.6662 0.7716 53.11 48.75 93.48 36.01
10 0.5399 0.5413 0.4058 0.6664 164.89 45.93 68.55 112.75
Comparison of the Selection of Variance